# Aperiodic Structures Never Collapse: Fibonacci Hierarchies for Lossless Compression

### 16 Mar 2026 v1

@@ 4.10 Theorem: Sturmian Codebook Efficiency
The Morse–Hedlund theorem [3] states that an infinite binary sequence is aperiodic if and only if $p(n)\geq n+1$ for all $n\geq 1$, and Sturmian if and only if equality holds throughout. The Fibonacci word is the canonical Sturmian sequence over $\{L,S\}$ (see also [11]), so $p(n)=n+1$ for all $n$. The supertile sequence at each hierarchy level $k$ is the image of the original Fibonacci word under $(\sigma^{-1})^{k}$; since $\sigma^{-1}$ preserves the PV fixed-point structure (Section 4.6), every level is also Sturmian and satisfies $p(n)=n+1$. A codebook of $C_{m}$ entries covers $\min(C_{m},F_{m}+1)$ of the $F_{m}+1$ distinct $F_{m}$-letter tile words, giving (17). Any aperiodic non-Sturmian sequence has $p(F_{m})\geq F_{m}+2$, reducing coverage to at most $C_{m}/(F_{m}+2)$. ∎
@@ 4.12 Theorem: Golden Compensation
since $(\psi/\varphi)^{m}\to 0$. For all finite $m$ the exact value is $C(m)=W(\varphi^{m}-\psi^{m})/(\sqrt{5}\,\varphi^{m-1})$. ∎
@@ 4.13 Theorem: Level Activation Threshold
The expected number of $F_{m}$-gram hits is $P(m)\cdot r_{m}=W\,r_{m}/\varphi^{m-1}$. Setting this $\geq T_{m}$ and solving for $W$ gives (26). ∎
@@
The threshold for at least one hit is easily exceeded; the more practically relevant threshold is when the codebook accumulates enough repeated $F_{m}$-grams to pass pruning and yield non-trivial entries. Empirically, the 89-gram and 144-gram codebooks are empty at enwik8 (W=27,700, [^0] W=27{,}700{,}000, 0 hits) and active at enwik9 (W=298,263,298W=298{,}263{,}298, giving 2,396 and 945 hits respectively), consistent with the codebook activation scaling as $O(\varphi^{m-1})$.
@@ 4.15 Theorem: Aperiodic Parsing Dominance
Since $H_{b}(r)\leq 1$ for all $r\in[0,1]$:
@@
using $\varphi^{2}=\varphi+1$ in the final step. The series converges because position counts decay geometrically with ratio $1/\varphi<1$. ∎
@@ 4.16 Theorem: Strict Coding Entropy for Long-Range Sources
Conditioning on additional context never increases entropy: for $k<k^{\prime}$,
@@
with strict inequality iff $X_{-k:-(k^{\prime}-1)}$ carries additional information about $X_{0}$; so $h_{k}$ is non-increasing in $k$.
Left inequality.  The Fibonacci tiling at level $m$ provides super-L positions with $F_{m}$-gram spans. A codebook approximating $P(X_{0}\mid X_{-1:-(F_{m}-1)})$ drives the per-word coding cost at those positions to $h_{F_{m}}(P)$, so $H^{\text{fib}}(P)\leq h_{F_{m}}(P)$.
Strict central inequality.  By $m$-LRPD: $h_{F_{m}}(P)<h_{F_{m-1}}(P)$. Since $m>m^{*}$, we have $F_{m-1}\geq F_{m^{*}}$, so $h_{F_{m-1}}(P)\leq h_{F_{m^{*}}}(P)$ by (35). Combining: $h_{F_{m}}(P)<h_{F_{m^{*}}}(P)$ strictly.
Right inequality.  The periodic tiling collapses at $m^{*}$: no super-L positions exist beyond depth $m^{*}$, so the coder is limited to contexts of length $\leq F_{m^{*}}$, giving $H^{\text{per}}(P)\geq h_{F_{m^{*}}}(P)$. ∎
@@ 4.17 Compression Bounds: Redundancy and Dictionary Efficiency
Since $h_{k}(P)$ is non-increasing and $h_{\infty}(P)=h(P)$:
$R_{k}(P)$ $=\sum_{n=k}^{\infty}\bigl[h_{n}(P)-h_{n+1}(P)\bigr]$
$=\sum_{n=k}^{\infty}I\!\left(X_{0};\,X_{-n}\mid X_{-1},\ldots,X_{-n+1}\right)$
$\leq\sum_{n=k}^{\infty}I(X_{0};\,X_{-n})$
$\leq C\sum_{n=k}^{\infty}e^{-n/\lambda}\;=\;\frac{C\,e^{-k/\lambda}}{1-e^{-1/\lambda}},$
using the data processing inequality in the second step and (36) in the third. Setting $k=F_{m}$ and using $F_{m}=\varphi^{m}/\sqrt{5}+O(|\psi|^{m})$ gives (38). The lower bound (39) holds for any $m$-LRPD source at scale $k=F_{m^{*}}$, by Definition 5. ∎
@@
since $C_{k_{\max}}\geq 1$. The periodic bound $\mathcal{V}^{\text{per}}=O(\varphi^{m^{*}})$ follows from terminating the sum at $m^{*}$ and bounding $F_{m}\leq F_{m^{*}}$ for $m\leq m^{*}$. ∎

### 23 Mar 2026 v2

@@ 4.10 Theorem: Sturmian Codebook Efficiency
Step 1: Why exactly $n+1$ factors (the Morse–Hedlund theorem).  The Morse–Hedlund theorem [3] states that an infinite binary sequence is aperiodic if and only if $p(n)\geq n+1$ for all $n\geq 1$, and Sturmian if and only if equality holds throughout.
The intuition is a sliding-window argument. Consider a window of length $n$ sliding along an infinite binary sequence. A periodic sequence of period $p$ can produce at most $p$ distinct windows (the window content repeats after $p$ shifts), so $p(n)\leq p$ — eventually $p(n)<n+1$ for $n\geq p$, which is why periodic sequences fail. An aperiodic sequence must keep producing new factors as $n$ grows, giving $p(n)\geq n+1$. The Sturmian sequences achieve this lower bound exactly: at each length $n$, exactly one “special” factor can be extended in two ways (by appending either $L$ or $S$), producing $p(n+1)=p(n)+1=(n+1)+1=n+2$. This is the slowest possible growth of factor complexity consistent with aperiodicity.
Step 2: Concrete example for $n=3$.  The Fibonacci word begins $u=L\,S\,L\,L\,S\,L\,S\,L\,L\,S\,L\,L\,S\,\ldots$ Sliding a window of length $n=3$, we find exactly $3+1=4$ distinct factors:
Factor First occurrence
LSLLSL positions 1–3
SLLSLL positions 2–4
LLSLLS positions 3–5
SLSSLS positions 5–7
No fifth factor of length 3 ever appears, no matter how far we extend the Fibonacci word. By contrast, a generic aperiodic binary sequence could have up to $2^{3}=8$ factors of length 3; the Fibonacci word has exactly 4, the minimum possible. The “missing” factors ($LLL$, SSSSSS, SSLSSL, LSSLSS) are forbidden by the balanced structure: the Sturmian property ensures that any two factors of equal length differ in their $L$-count by at most 1.
Step 3: Sturmianity is preserved under deflation.  The Fibonacci word is the canonical Sturmian sequence over $\{L,S\}$ (see also [11]), so $p(n)=n+1$ for all $n$. The supertile sequence at each hierarchy level $k$ is the image of the original Fibonacci word under $(\sigma^{-1})^{k}$; since $\sigma^{-1}$ preserves the PV fixed-point structure (Section 4.6), every level is also Sturmian and satisfies $p(n)=n+1$.
Why does deflation preserve Sturmianity? The key is that the Fibonacci substitution $\sigma$ is a morphism of Sturmian sequences: if $u$ is the Sturmian sequence with slope $1/\varphi$, then $\sigma^{-1}(u)$ is the Sturmian sequence with slope $1/\varphi^{2}$ (still irrational, since $\varphi$ is a PV number). More generally, $(\sigma^{-1})^{k}(u)$ is Sturmian with slope $1/\varphi^{k+1}$, which is irrational for all finite $k$. Since every irrational rotation coding is Sturmian, the factor complexity $p(n)=n+1$ holds at every hierarchy level.
Step 4: Codebook efficiency bound.  A codebook of $C_{m}$ entries covers $\min(C_{m},F_{m}+1)$ of the $F_{m}+1$ distinct $F_{m}$-letter tile words, giving (17). Any aperiodic non-Sturmian sequence has $p(F_{m})\geq F_{m}+2$, reducing coverage to at most $C_{m}/(F_{m}+2)$. The efficiency gap $C_{m}/(F_{m}+1)-C_{m}/(F_{m}+2)=C_{m}/[(F_{m}+1)(F_{m}+2)]$ is small per level but cumulates across all $k_{\max}$ levels of the hierarchy. ∎
@@ 4.12 Theorem: Golden Compensation
since $(\psi/\varphi)^{m}\to 0$. For all finite $m$ the exact value is $C(m)=W(\varphi^{m}-\psi^{m})/(\sqrt{5}\,\varphi^{m-1})$.
Numerical verification.  The following table evaluates $C(m)=P(m)\cdot F_{m}=W\cdot F_{m}/\varphi^{m-1}$ for each hierarchy level, using W=298,263,298W=298{,}263{,}298 (enwik9). The limit is $W\varphi/\sqrt{5}=298{,}263{,}298\times 0.72361\approx 215{,}841{,}135$.
$m$ $F_{m}$ $P(m)$ $C(m){=}P{\cdot}F_{m}$ Dev.
4 3 70.4M 211.2M $2.1\%$
5 5 43.5M 217.6M $0.8\%$
6 8 26.9M 215.2M $0.3\%$
7 13 16.6M 216.0M $0.1\%$
8 21 10.3M 215.7M $0.05\%$
9 34 6.3M 215.8M $0.004\%$
10 55 3.9M 215.8M $0.008\%$
11 89 2.4M 215.8M $0.005\%$
12 144 1.5M 215.8M $0.007\%$
W=298,263,298W=298{,}263{,}298. Limit: $W\varphi/\sqrt{5}\approx 215.8$M.
The convergence is rapid: $C(m)$ oscillates around the limit $\approx 215{,}841{,}135$ with deviations of less than 2.2% already at $m=4$ and less than 0.01% by $m=9$. The alternating sign of the deviation comes from the $\psi^{m}$ correction term, which alternates in sign because $\psi<0$. ∎
@@ 4.13 Theorem: Level Activation Threshold
The argument proceeds in three steps.
Step 1: Count the available positions.  At the $F_{m}$-gram level, the number of super-L positions in a Fibonacci tiling of $W$ words is $P(m)=W/\varphi^{m-1}$ (Eq. (23)). Each such position is a candidate for an $F_{m}$-gram codebook lookup.
Step 2: Apply the hit rate.  Not every candidate position produces a codebook hit — only those whose $F_{m}$-word phrase matches an entry in the codebook. The fraction that matches is the hit rate $r_{m}$. So the expected number of $F_{m}$-gram hits is:
$\text{Expected hits}=P(m)\cdot r_{m}=\frac{W\cdot r_{m}}{\varphi^{m-1}}.$
Step 3: Solve for the threshold.  For the level to contribute at least $T_{m}$ hits, we need:
$\frac{W\cdot r_{m}}{\varphi^{m-1}}\geq T_{m}\;\Longrightarrow\;W\geq\frac{T_{m}\cdot\varphi^{m-1}}{r_{m}}=:W^{*}_{m}.$
This is Eq. (26).
Concrete computation for 89-gram and 144-gram with enwik9.  For the 89-gram level ($m=11$): $\varphi^{10}\approx 122.99$, and the empirical hit rate at enwik9 is $r_{11}\approx 0.001$ (0.10%). The expected number of hits at W=298,263,298W=298{,}263{,}298 is:
$P(11)\cdot r_{11}=\frac{298{,}263{,}298}{122.99}\times 0.001\approx 2{,}425\;\text{hits},$
which matches the empirical count of 2,3962{,}396. The minimum threshold for a single hit is $W^{*}_{11}=1\times 122.99/0.001\approx 123{,}000$ words — easily exceeded by enwik9.
For the 144-gram level ($m=12$): $\varphi^{11}\approx 199.01$, $r_{12}\approx 0.0006$ (0.06%). Expected hits:
$P(12)\cdot r_{12}=\frac{298{,}263{,}298}{199.01}\times 0.0006\approx 899\;\text{hits},$
consistent with the empirical count of $945$. The threshold is $W^{*}_{12}=199.01/0.0006\approx 332{,}000$ words.
At enwik8 (W=27,731,124W=27{,}731{,}124), the expected hit counts are $27{,}731{,}124/122.99\times 0.001\approx 225$ (89-gram) and $27{,}731{,}124/199.01\times 0.0006\approx 84$ (144-gram). However, both levels show 0 empirical hits at enwik8 because the codebook requires repeated $F_{m}$-grams to build entries: the effective $T_{m}$ for non-trivial codebook formation is much larger than 1, as discussed below. ∎
@@
The threshold for at least one hit is easily exceeded; the more practically relevant threshold is when the codebook accumulates enough repeated $F_{m}$-grams to pass pruning and yield non-trivial entries. The effective $T_{m}$ for codebook formation is on the order of hundreds to thousands, because a phrase must appear multiple times across the corpus before it is promoted to a codebook entry. Empirically, the 89-gram and 144-gram codebooks are empty at enwik8 (W=27,700, [^0] W=27{,}700{,}000, 0 hits) and active at enwik9 (W=298,263,298W=298{,}263{,}298, giving 2,396 and 945 hits respectively), consistent with the codebook activation scaling as $O(\varphi^{m-1})$.
@@ 4.15 Theorem: Aperiodic Parsing Dominance
Step 1: Upper-bound each term.  The binary entropy function satisfies $H_{b}(r)\leq 1$ for all $r\in[0,1]$, with equality at $r=1/2$. At each hierarchy level $k$, the number of positions that require a flag bit is $P(k)=W/[\varphi^{k+1}(\varphi+1)]$ (from Eq. (18)). The per-word flag entropy (Eq. (29)) can therefore be bounded term by term:
$\frac{P(k)\cdot H_{b}(r_{k})}{W}\;\leq\;\frac{P(k)}{W}\;=\;\frac{1}{\varphi^{k+1}(\varphi+1)}.$
Step 2: Sum the geometric series.  From Step 1, $P(k)/W=1/[\varphi^{k+1}(\varphi+1)]$. Summing over all levels $k=1,2,3,\ldots$:
$h_{\text{flags}}^{\text{fib}}$ $\;\leq\;\sum_{k=1}^{\infty}\frac{1}{\varphi^{k+1}(\varphi+1)}\;=\;\frac{1}{\varphi(\varphi+1)}\sum_{k=1}^{\infty}\varphi^{-k}.$
The inner sum is a geometric series with common ratio $1/\varphi\approx 0.618<1$, so it converges. By the standard formula $\sum_{k=1}^{\infty}x^{k}=x/(1-x)$ with $x=1/\varphi$:
$\sum_{k=1}^{\infty}\varphi^{-k}=\frac{1/\varphi}{1-1/\varphi}=\frac{1}{\varphi-1}.$
Now we use the golden ratio identity $\varphi-1=1/\varphi$ (equivalent to $\varphi^{2}-\varphi-1=0$), giving $\sum_{k=1}^{\infty}\varphi^{-k}=\varphi$. Substituting back:
$h_{\text{flags}}^{\text{fib}}$ $\;\leq\;\frac{1}{\varphi(\varphi+1)}\cdot\varphi\;=\;\frac{1}{\varphi+1}\;=\;\frac{1}{\varphi^{2}}\;=\;\frac{1}{\varphi}\cdot\frac{1}{\varphi},$
where the second equality uses $\varphi+1=\varphi^{2}$. Equivalently, one can write this more directly as:
@@
using $\varphi^{2}-1=(\varphi+1)-1=\varphi$ in the final step. The series converges because position counts decay geometrically with ratio $1/\varphi<1$.
Step 3: Intuition — why the series converges.  The convergence of the flag overhead is a geometric series argument: level 1 contributes at most $1/[\varphi^{2}(\varphi+1)]\approx 0.146$ bits/word of flag overhead, level 2 contributes at most $1/[\varphi^{3}(\varphi+1)]\approx 0.090$ bits/word, level 3 at most $\approx 0.056$, and so on. Each successive level contributes $1/\varphi\approx 0.618$ times as much overhead as the previous level. The partial sums are:
Levels Cumul. (b/w) % limit
1 0.146 23.6
1–2 0.236 38.2
1–3 0.292 47.2
1–4 0.326 52.8
[…]
The bound $1/\varphi\approx 0.618$ bits/word is already tight after relatively few levels. Even with infinitely many active hierarchy levels, the total flag cost never exceeds $0.618$ bits per word.
Step 4: Practical meaning of $1/\varphi$ bits/word.  What does $1/\varphi\approx 0.618$ bits/word mean concretely? For enwik9 (W=298,263,298W=298{,}263{,}298 words), the worst-case total flag overhead across the entire hierarchy is bounded by $298{,}263{,}298\times 0.618/8\approx 23{,}028{,}000$ bytes. This is about 2.3% of the original 1 GB file — a small fixed price for access to codebook lookups at all hierarchy depths. In practice, the actual overhead is much smaller because hit rates $r_{k}$ are well below $1/2$, making $H_{b}(r_{k})\ll 1$. ∎
@@ 4.16 Theorem: Strict Coding Entropy for Long-Range Sources
The proof establishes a chain of four inequalities. We explain each link in the chain separately, then assemble them.
Background: conditioning reduces entropy.  Conditioning on additional context never increases entropy: for $k<k^{\prime}$,
@@
with strict inequality iff $X_{-k:-(k^{\prime}-1)}$ carries additional information about $X_{0}$; so $h_{k}$ is non-increasing in $k$. This is a direct consequence of the chain rule for entropy: knowing more context can only help (or be neutral for) predicting the next word.
Left inequality: $H^{\text{fib}}(P)\leq h_{F_{m}}(P)$.  The Fibonacci tiling at level $m$ provides super-L positions with $F_{m}$-gram spans. A codebook approximating $P(X_{0}\mid X_{-1:-(F_{m}-1)})$ drives the per-word coding cost at those positions to $h_{F_{m}}(P)$, so $H^{\text{fib}}(P)\leq h_{F_{m}}(P)$.
This is because the Fibonacci hierarchy provides context windows of length $F_{m}$ at level $m$ — a coder exploiting these windows achieves per-word entropy no worse than the $F_{m}$-gram conditional entropy. In practice, $H^{\text{fib}}(P)$ can be even lower than $h_{F_{m}}(P)$ because the Fibonacci hierarchy provides contexts at all levels simultaneously: $F_{4}=3$, $F_{5}=5$, …, $F_{m}$. The coder selects the best available context at each position.
Strict central inequality: $h_{F_{m}}(P)<h_{F_{m^{*}}}(P)$.  By $m$-LRPD (Definition 5): $h_{F_{m}}(P)<h_{F_{m-1}}(P)$. Since $m>m^{*}$, we have $F_{m-1}\geq F_{m^{*}}$, so $h_{F_{m-1}}(P)\leq h_{F_{m^{*}}}(P)$ by (35). Combining: $h_{F_{m}}(P)<h_{F_{m^{*}}}(P)$ strictly.
This is the key step: the $m$-LRPD condition says that context windows of $F_{m}$ words carry strictly more predictive information than windows of $F_{m-1}$ words. In natural language, this corresponds to the empirical observation that paragraph-level context (e.g., 89 or 144 words) genuinely helps predict the next word better than sentence-level context (e.g., 5 or 8 words). This is not surprising: Wikipedia articles maintain topical coherence across paragraphs, technical terms recur within sections, and stylistic patterns persist over hundreds of words. The $m$-LRPD condition is the formal expression of this long-range coherence.
Right inequality: $h_{F_{m^{*}}}(P)\leq H^{\text{per}}(P)$.  The periodic tiling collapses at $m^{*}$: no super-L positions exist beyond depth $m^{*}$, so the coder is limited to contexts of length $\leq F_{m^{*}}$, giving $H^{\text{per}}(P)\geq h_{F_{m^{*}}}(P)$.
In other words, the periodic tiling’s coder cannot exploit any context longer than $F_{m^{*}}$ words, because the hierarchy has collapsed and provides no positions for deeper lookups. Even if the source has useful long-range structure, the periodic coder cannot see it.
Assembly.  Chaining the three inequalities:
$\underbrace{H^{\text{fib}}(P)}_{\text{Fib.}}\;\leq\;\underbrace{h_{F_{m}}(P)}_{F_{m}\text{-gram}}\;\underbrace{<}_{m\text{-LRPD}}\;\underbrace{h_{F_{m^{*}}}(P)}_{F_{m^{*}}\text{-gram}}\;\leq\;\underbrace{H^{\text{per}}(P)}_{\text{per.}}$
The strict inequality in the middle makes the overall inequality strict: $H^{\text{fib}}(P)<H^{\text{per}}(P)$. ∎
@@ 4.17 Compression Bounds: Redundancy and Dictionary Efficiency
We proceed in four steps: telescoping the redundancy, applying the data processing inequality, bounding with the mixing condition, and specialising to Fibonacci indices.
Step 1: Telescope the redundancy.  Since $h_{k}(P)$ is non-increasing in $k$ (more context never hurts) and $h_{\infty}(P)=h(P)$ is the true entropy rate, the redundancy is the total “room for improvement” beyond context length $k$:
$R_{k}(P)$ $=h_{k}(P)-h(P)=\sum_{n=k}^{\infty}\bigl[h_{n}(P)-h_{n+1}(P)\bigr].$
Each term $h_{n}(P)-h_{n+1}(P)$ is the entropy reduction gained by extending the context window from $n$ to $n+1$ words. By the chain rule for conditional entropy, this equals the conditional mutual information:
$h_{n}(P)-h_{n+1}(P)=I\!\left(X_{0};\,X_{-n}\mid X_{-1},\ldots,X_{-n+1}\right).$
This is the amount of new information that the single extra word $X_{-n}$ provides about $X_{0}$, beyond what $X_{-1},\ldots,X_{-n+1}$ already tell us.
Step 2: Apply the data processing inequality.  The data processing inequality states that conditioning on additional variables (the intermediate words $X_{-1},\ldots,X_{-n+1}$) can only reduce mutual information:
$I\!\left(X_{0};\,X_{-n}\mid X_{-1},\ldots,X_{-n+1}\right)\;\leq\;I(X_{0};\,X_{-n}).$
Intuitively: if $X_{-n}$ tells you at most $I(X_{0};X_{-n})$ bits about $X_{0}$ when you have no other context, it cannot tell you more than that when you already know the intermediate words. This gives:
$R_{k}(P)$ $=\sum_{n=k}^{\infty}I\!\left(X_{0};\,X_{-n}\mid X_{-1},\ldots,X_{-n+1}\right)$
$\leq\sum_{n=k}^{\infty}I(X_{0};\,X_{-n}).$
Step 3: Apply the exponential mixing bound.  By the mixing assumption (Definition 6), $I(X_{0};X_{-n})\leq C\,e^{-n/\lambda}$. Substituting:
$R_{k}(P)$ $\leq C\!\sum_{n=k}^{\infty}e^{-n/\lambda}=C\,e^{-k/\lambda}\!\sum_{j=0}^{\infty}e^{-j/\lambda}$
$=\frac{C\,e^{-k/\lambda}}{1-e^{-1/\lambda}},$
using the geometric series ($e^{-1/\lambda}<1$).
Step 4: Specialise to Fibonacci indices.  Setting $k=F_{m}$ and using $F_{m}=\varphi^{m}/\sqrt{5}+O(|\psi|^{m})$ gives (38). The lower bound (39) holds for any $m$-LRPD source at scale $k=F_{m^{*}}$, by Definition 5.
Numerical example with $\lambda\approx 20$.  For natural language, a reasonable estimate of the mixing scale is $\lambda\approx 20$ words (mutual information between words decays to negligible levels beyond about 20 positions). Using $C=1$ bit for simplicity, and the normalisation $1/(1-e^{-1/\lambda})=1/(1-e^{-1/20})\approx 20.5$:
$m$ $F_{m}$ $R_{F_{m}}$ bound (bits) Ratio
4 3 17.5 1
5 5 15.9 0.91
6 8 13.7 0.78
7 13 10.7 0.61
8 21 7.2 0.41
9 34 3.7 0.21
10 55 1.3 0.075
11 89 0.024 0.0014
12 144 0.002 $9{\times}10^{-5}$
The redundancy drops from $17.5$ bits at $F_{4}=3$ to $0.024$ bits at $F_{11}=89$ — a $700\times$ reduction. At $F_{12}=144$, the redundancy is effectively zero (<0.002<0.002 bits). A periodic tiling that collapses at $m^{*}=4$ ($F_{4}=3$) is permanently stuck at $R_{3}(P)\leq 17.5$ bits of residual redundancy per word, while the Fibonacci hierarchy drives this to negligible levels by level 11. ∎
@@
since $C_{k_{\max}}\geq 1$. The periodic bound $\mathcal{V}^{\text{per}}=O(\varphi^{m^{*}})$ follows from terminating the sum at $m^{*}$ and bounding $F_{m}\leq F_{m^{*}}$ for $m\leq m^{*}$.
Numerical table of $E_{m}$ for $m=1,\ldots,9$.  Using $\bar{h}\approx 5$ bits/word (a representative per-word entropy for English text under arithmetic coding) and empirical codebook sizes from enwik9:
$m$ $F_{m}$ $C_{m}$ $F_{m}\bar{h}$ $E_{m}$
1 2 8K 10 3.0-3.0
2 3 8K 15 2.0
3 5 6K 25 12.4
4 8 4K 40 28.0
5 13 2.5K 65 53.7
6 21 1.5K 105 94.5
7 34 800 170 160.4
8 55 300 275 266.8
9 89 100 445 438.4
$E_{m}=F_{m}\bar{h}-\log_{2}C_{m}$ with $\bar{h}\approx 5$ bits/word.
The per-entry gain $E_{m}$ grows roughly as $5\cdot F_{m}$ (the $\log_{2}C_{m}$ overhead becomes negligible for large $m$). At the 89-gram level ($m=9$), each codebook hit saves approximately $438$ bits — it encodes 89 words with a single codebook reference instead of 89 individual arithmetic coding symbols. The codebook address costs only $\log_{2}(100)\approx 6.6$ bits. The ratio $E_{9}/E_{2}=438.4/2.0=219$ reflects the exponential growth.
Note that $E_{1}<0$: at the bigram level, codebook overhead exceeds the savings because $\log_{2}C_{1}=13$ bits is larger than the $F_{1}\bar{h}=10$ bits saved. This is consistent with the observation that bigram codebook hits are only useful when the codebook is small enough (i.e., when phrase repetition is frequent enough to justify the addressing cost). At deeper levels, $F_{m}\bar{h}$ dominates and every hit is strongly net-positive. ∎